Logiweb diagnose of opgave

Up Help

No errors found

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-30.UTC:01:42:49.465331 = MJD-53916.TAI:01:43:22.465331 = LGT-4658348602465331e-6